OpenStudy (anonymous):

What is the radius of the circle whose equation is x^2 + y^2 = 9? 9?

OpenStudy (anonymous):

3

OpenStudy (lgbasallote):

nope...here's a hint the equation of a circle is (x-h)^2 + (y-k)^2 = r^2

OpenStudy (lgbasallote):

so 9 is just r^2 to find r you need to take the square root..make sense?

OpenStudy (anonymous):

A cricle centered at the origin has the equation \[x^2 + y^2 = r^2\]if \(r^2\) is 9, what is r, the radius?

OpenStudy (anonymous):